Company X has a beta of 1.75, while Company Y has a beta of 2.1. The return on the market is 14%, and the risk-free rate is 7%. By how much does Company’s required return exceed Company’s required return?
1
Expert's answer
2020-05-11T07:41:06-0400
"ri=rRF+(rM-rRF)\times (\beta)"

"ri_1=7+(14-7)\times (1.75)=19.25\%"

"ri_2=7+(14-7)\times (2.1)=21.7\%"

"21.7\%-19.25\%=2.45\%"

 Company’s Y required return exceeds Company’s X required return by "2.45\%."